# Repository-level Context Code Translation Benchmark Targeting Rust

### v5

@@ 4.5.2. Results
Missing Punctuation Marks. This error cause arises from the LLM’s failure to include necessary punctuation in its generated output. Examples include incomplete parentheses, missing commas between parameters, and other critical punctuation omissions.
@@
Overall Distribution of Error Causes. As shown in Fig. 7, the most common error cause encountered by LLMs during the code translation task on RustRepoTrans is Failing to Understand Differences Between Languages, which accounts for 77.9% of the errors. This is followed by Failing to Understand Target Language Features at 18.5%, and Others at 3.7%. These results indicate that the most significant challenge for LLMs in code translation is effectively grasping the various differences between the source and target languages, such as dependency and syntax differences.
Figure 7. Overall Distribution of Error Causes
At a more detailed level, the top three error causes are Function Differences Misinterpretation (36.6%), Variable Differences Misinterpretation (25.3%), and Data Type Misinterpretation (14.6%). This distribution highlights that function and variable dependencies constitute the largest portion of the dependencies involved in RustRepoTrans code translation tasks. Furthermore, since RustRepoTrans targets a strongly typed language, errors related to data types are ranked third.
Figure 8. Comparative Distribution of Error Causes in LLMs
Comparison of Error Causes Distribution across Different LLMs. Fig. 8 shows the distribution of error causes across various LLMs during code translation on RustRepoTrans. The Failing to Understand Target Language Features error causes are relatively evenly distributed among LLMs due to their similar training data, primarily derived from publicly available code. This results in comparable understanding of the target language across models. In contrast, Failing to Understand Differences Between Languages and Other errors are more common in LLMs with weaker translation capabilities. The former evaluates the model’s understanding of language differences, while the latter reflects its ability to follow instructions, both positively correlating with translation performance.
Notably, Claude-3.5, the top-performing model on RustRepoTrans, exhibits the lowest proportions of Failing to Understand Differences Between Languages and Other error causes. For simpler error types like Syntactic Differences Misinterpretation, Dependency Resolution Differences Misinterpretation, Missing Punctuation Marks, and Function Signature Inconsistencies, the proportions were as low as 2.8%, 2.5%, 0%, and 0%, respectively, indicating superior understanding and adherence to instructions compared to other models.
4.5.3. Conclusion
@@
4.6. RQ5: Key Capabilities
@@ 4.6.1. Design
The key abilities for these three aspects were tested by evaluating each model on a dataset extracted or constructed from RustRepoTrans, tailored to the specific capability being assessed. The same translation prompts used in Section 4.2 were applied. The experimental design for each aspect is described as follows.
@@
Noise Robustness. This capability evaluates the model’s capacity to identify necessary dependencies from provided options, which is crucial in real-world scenarios where dependency information is often uncertain or incomplete. We assessed this capability through two angles: redundancy and incompleteness. For redundancy, we created a dataset by selecting functions and data types with high text similarity (using BLEU scores (Papineni et al., 2002)) relative to the target function’s dependencies, excluding those in the original set. We introduce a novel metric, Redundancy Impact Rate (RIR), to measure the ratio of successful translations between scenarios with Redundant Dependencies and All Dependencies. For incompleteness, we created various datasets by randomly reducing the target function’s dependencies to 75%, 50%, 25%, and 0% of their original size. Using these new benchmarks, we assessed the success rates (Pass@1) of the LLMs and compared them to the original Pass@1 with all dependencies (see RQ1). We introduce a novel metric, Incompleteness Impact Rate (IIR), which measures the LLM’s performance by calculating the average of the Pass@1 under 100%(i.e. All), 75%, 50%, 25%, and 0% dependencies.
@@
Syntactical Differences Identification. This capability assesses the model’s skill in recognizing syntactical differences between source and target programming languages, a key requirement for accurate code translation. For instance, Rust, as a strongly-typed language, does not require checks such as memory exception handling (common in C), type checks (in Python), or null pointer checks (in C and Java) shown in Fig. 9. We selected 89 function pairs from RustRepoTrans that include these checks in the source language and evaluated whether the LLM correctly omits them in the target language. We introduce a novel metric, Syntactical Differences Identification Rate (SDIR@K), which measures the LLM’s success in identifying and omitting these checks in the translated code across $K$ samples.
@@
Figure 9. Examples of checks from other languages that are not needed in Rust
@@
Code Simplicity. This capability assesses the simplicity of translation results, as simpler code is preferred by developers. We evaluate this by calculating token counts with tree-sitter (tree sitter, 2023) and measuring cyclomatic complexity (Gill and Kemerer, 1991), where higher values indicate greater complexity. We apply rust-code-analysis (Ardito et al., 2020) to both the reference and translated code that passed the test cases. A higher ratio suggests that the translated code is more concise and closely matches the reference. We introduce two metrics for code simplicity: Token Rate and CC Rate. These compare the token counts and cyclomatic complexities of the reference and translated code. Higher values for both metrics indicate that the generated code is simpler and more aligned with the original.
Figure 10. The Pass@1 of LLMs under different proportions of dependencies
4.6.2. Results
@@
Noise Robustness. From the perspective of redundant dependencies, after introducing interference dependencies, Claude-3.5 and DeepSeekCoderV2-16B demonstrated minimal fluctuations, with RIR values of 99.4% and 101.6%. These variations fall within the expected range for LLM generation. In contrast, GPT-4 and Llama-3.1-8B exhibited larger fluctuations, with RIR values of 88.2% and 92%, highlighting potential areas for improvement in their noise robustness regarding redundant dependencies. Regarding incomplete dependencies, as shown in Fig. 10, DeepSeekCoderV2-16B and Llama-3.1-8B showed stable performance during gradual reductions, with drop ratios of 19.1% and 18.0% at the 25% level. Noticeable declines occurred only at the Without Dependencies stage. In contrast, Claude-3.5 and GPT-4 faced significant drops at nearly every stage, likely because their successful translations often rely on multiple dependencies. Most successful translations for DeepSeekCoderV2-16B and Llama-3.1-8B came from simpler data with fewer dependencies, ensuring consistent availability across stages except at the Without Dependencies stage. As shown in Fig. 11, the IIR ranking among models are the same as the Pass@1 under All Dependencies in 4.2, Claude-3.5 demonstrated the best performance with IIR values of 32.2%.
@@
Syntactical Differences Identification. As shown in Fig. 11, Claude-3.5 achieves the highest ISDR@1 at 95.5%, demonstrating a strong ability to identify syntactical differences between programming languages. This finding is consistent with the results in Section 4.4 regarding the distribution of error causes across different LLMs (see Fig. 8). In contrast, DeepSeekCoderV2-16B has the lowest ISDR@1 at only 40.5%. This indicates that DeepSeekCoderV2-16B often assumes the checks performed in the source language are also necessary in the target language, leading to errors by incorrectly treating those checks as valid in the translation.
@@
Code Simplicity. Fig. 11 shows that among the successfully translated tasks, the Token Rate for code generated by Claude-3.5 and GPT-4 reached 96.6% and 97.0%, respectively, indicating high simplicity compared to the reference code. In contrast, Llama-3.1-8B had the lowest Token Rate at 63.1%, suggesting it requires more code to achieve the same functionality.
@@
In terms of CC Rate, Claude-3.5 again excels with a CC Rate of 103.7%, while Llama-3.1-8B is notably lower at 53.6%. This difference arises because Claude-3.5 often uses concise built-in functions in place of loops. For instance, as illustrated in Fig. 12, when tasked with finding the first element in an array that meets a condition, Claude-3.5 employs the target language’s any method from the iterator library, whereas other LLMs rely on traditional loops. Furthermore, Claude-3.5’s CC Rate exceeding 1 indicates that, in terms of cyclomatic complexity, its generated code is simpler than the reference code. Fig. 12 shows an example with C as the source language. This simplification likely results from Rust ground truth developers using the C2Rust migration tool (C2Rust, 2024), which translates code through one-to-one pattern matching.
@@
These findings suggest that Claude-3.5 demonstrates a stronger code comprehension and implementation ability, accurately understanding the source code’s functionality and translating it concisely based on the target language’s features.
Figure 11. The performance of LLMs on Key Abilities of code translation.
@@
Figure 12. Claude-3.5 translation result compared to Other LLMs and reference code
4.6.3. Conclusion
@@
The complexity of code generated by LLMs in translation reflects their translation capabilities: models with stronger abilities produce code with lower complexity for equivalent functionality. Claude-3.5, which excels on RustRepoTrans, also shows superior performance in Noise Robustness, Syntactical Differences Identification, and Code Simplicity.
5. Threats to Validity
One potential threat is data leakage between our benchmark and model training data. However, the training data comprises independent function code from different languages rather than functionally equivalent code pairs, which is crucial for code translation. Another concern is the limited size and variety of programming languages in our benchmark, which may impact the generalizability of our findings. We plan to extend our benchmark in the future. Some projects also lack consistent documentation across language versions, leading to potential functional discrepancies. We mitigated this by manually verifying the functionality of the final function pairs. Additionally, the static code analysis tool tree-sitter may not accurately retrieve all dependencies for Rust functions, which we addressed through manual validation. Lastly, our study involved only four models due to time and resource constraints. The focus of this paper is the benchmark itself, not the models. We intend to incorporate more models in future work, but the conclusions and evaluation capabilities of our benchmark remain valid.
6. Conclusion
@@
This work makes the first attempt to evaluate LLMs on repository-level code generation targeting Rust. We first manually construct the first repository-level code translation benchmark RustRepoTrans and perform the first study of 4 state-of-the-art LLMs on repository-level code translation. we find that existing LLMs show much worse performance on repository-level code translation compared to on standalone code translation, LLMs still have significant room for improvement in their code translation capabilities within real-world software development scenarios. Besides, when the target language is a low-resource language with multiple syntactic constraints, such as Rust, LLMs struggle to effectively identify the various differences between the source and target languages, as well as understand the features of the target language.

### v6

@@ IV-D2 Results
Missing Punctuation Marks. This error cause arises from the LLM’s failure to include necessary punctuation in its output. Examples include incomplete parentheses, missing commas between parameters, and other critical punctuation omissions.
@@
Overall Distribution of Error Causes. As shown in Fig. 7, the most common error cause encountered by LLMs during the code translation task on RustRepoTrans is Failing to Understand Differences Between Languages, which accounts for 73.9% of the errors. This is followed by Failing to Understand Target Language Features at 22.4%, and Others at 3.7%. These results indicate that the most significant challenge for LLMs in code translation is effectively grasping the various differences between the source and target languages, such as dependency and syntax differences.
At a more detailed level, the top three error causes are Function Differences Misinterpretation (38.6%), Variable Differences Misinterpretation (24.9%), and Data Type Misinterpretation (16.1%). This distribution highlights that function and variable dependencies constitute the largest portion of the dependencies involved in RustRepoTrans code translation tasks. Furthermore, since RustRepoTrans targets a strongly typed language, errors related to data types are ranked third.
Comparison of Error Causes Distribution across Different LLMs. Fig. 8 shows the distribution of error causes across various LLMs during code translation on RustRepoTrans. The Failing to Understand Target Language Features error causes are relatively evenly distributed among LLMs due to their similar training data, primarily derived publicly. This results in comparable understanding of the target language across models. In contrast, Failing to Understand Differences Between Languages and Other errors are more common in LLMs with weaker translation capabilities. The former evaluates the model’s understanding of language differences, while the latter reflects its ability to follow instructions, both positively correlating with translation performance.
Notably, DeepSeek-R1, the top-performing model on RustRepoTrans, exhibits the lowest rate of Syntactic, Function and Variable Differences Misinterpretation, representing its exceptional comprehension capabilities in capturing cross-lingual discrepancies. Furthermore, Claude-3.5 made no errors on simpler error types such as Missing Punctuation Marks and Function Signature Inconsistencies, indicating superior understanding and adherence to instructions compared to others.
IV-D3 Conclusion
@@
Figure 7: Overall Distribution of Error Causes
Figure 8: Comparative Distribution of Error Causes in LLMs
IV-E RQ4: Key Capabilities
@@ IV-E1 Design
The key abilities for these three aspects were tested by evaluating each model on a dataset extracted or constructed from RustRepoTrans, tailored to the specific capability being assessed. The same translation prompts used in Section IV-B were applied. The experimental design for each aspect is described as follows.
@@
Noise Robustness. This capability evaluates the model’s capacity to identify necessary dependencies from provided options, which is crucial in real-world scenarios where dependency information is often uncertain or incomplete. We assessed this capability through two angles: redundancy and incompleteness. For redundancy, we created a dataset by selecting functions and data types with high text similarity (using BLEU scores [63]) relative to the target function’s dependencies, excluding those in the original set. We introduce a novel metric, Redundancy Impact Rate (RIR), to measure the ratio of successful translations between scenarios with Redundant Dependencies and All Dependencies. For incompleteness, we created various datasets by randomly reducing the target function’s dependencies to 75%, 50%, 25%, and 0%. Using these new benchmarks, we assessed the success rates (Pass@1) of the LLMs and compared them to the original Pass@1 with all dependencies. We introduce a novel metric, Incompleteness Impact Rate (IIR), which measures the LLM’s performance by calculating the average of the Pass@1 under 100%(i.e. All), 75%, 50%, 25%, and 0% dependencies.
@@
Syntactical Differences Identification. This capability assesses the model’s skill in recognizing syntactical differences between source and target languages, a key requirement for accurate code translation. For instance, Rust, as a strongly-typed language, does not require checks such as memory exception handling (common in C), type checks (in Python), or null pointer checks (in C and Java) shown in Fig. 9. We selected 89 function pairs from RustRepoTrans that include these checks in the source language and evaluated whether the LLM correctly identifies and omits in the target language. We introduce a novel metric, Syntactical Differences Identification Rate (SDIR@K), which measures the LLM’s success in identifying these syntactical differences across $K$ samples.
@@
Figure 9: Examples of checks from other languages that are not needed in Rust
@@
Code Simplicity. This capability assesses the simplicity of translation results, as simpler code is preferred by developers. We evaluate this by calculating token counts with tree-sitter [40] and measuring cyclomatic complexity [64], where higher values indicate greater complexity. We apply rust-code-analysis [65] to the reference and translated code that passed the test cases. A higher ratio suggests that the translated code is more concise and closely matches the reference. We introduce two metrics for code simplicity: Token Rate and CC Rate. These compare the token counts and cyclomatic complexities of the reference and translated code. Higher values for both metrics indicate simpler generated code and is more aligned with the original.
Figure 10: The Pass@1 of LLMs under different proportions of dependencies
IV-E2 Results
@@
Noise Robustness. From the perspective of redundant dependencies, after introducing interference dependencies, Claude-3.5 and DeepSeekCoderV2-16B demonstrated minimal fluctuations, with RIR values of 99.4% and 101.6%. These variations fall within the expected range for LLM generation. In contrast, DeepSeek-R1 and DeepSeek-V3, the two LLMs that ranked top-2 in terms of Pass@1 and DSR@1, exhibited larger fluctuations, with RIR values of 92.7% and 86.2%, highlighting potential areas for improvement in their noise robustness regarding redundant dependencies. From the perspective of incomplete dependencies, as shown in Fig. 10, DeepSeekCoderV2-16B and Llama-3.1-8B showed stable performance during gradual reductions, with drop ratios of 19.1% and 18.0% at the 25% level. Noticeable declines occurred only at the Without Dependencies stage. This is because most successful translations for DeepSeekCoderV2-16B and Llama-3.1-8B came from simpler data with fewer dependencies, ensuring consistent availability across stages except at the Without Dependencies stage. In contrast, other LLMs faced significant drops at nearly every stage, likely because their successfully translated tasks include those with multiple dependencies. As shown in Fig. 11, the IIR ranking among models are almost the same as the Pass@1 under All Dependencies in IV-B, DeepSeek-R1 demonstrated the best performance with IIR values of 35.8%.
@@
Syntactical Differences Identification. As shown in Fig. 11, DeepSeek-V3, DeepSeek-R1 and Claude-3.5 achieve high SDIR@1 at 97.8%, 96.6% and 95.5% separately, demonstrating a strong ability to identify syntactical differences between programming languages. This finding is consistent with the results in Section IV-D regarding the distribution of error causes across different LLMs (see Fig. 8). In contrast, DeepSeekCoderV2-16B has the lowest SDIR@1 at only 40.5%. This indicates that DeepSeekCoderV2-16B often assumes the checks performed in the source language are also necessary in the target language, leading to errors by incorrectly treating those checks as valid in the translation.
@@
Code Simplicity. Fig. 11 shows that among the successfully translated tasks, the Token Rate for code generated by DeepSeek-R1 and DeepSeek-V3 reached 100.9% and 99.5%, respectively, indicating high simplicity compared to the reference code. In contrast, Llama-3.1-8B had the lowest Token Rate at 63.1%, suggesting it requires more code to achieve the same functionality.
@@
In terms of CC Rate, Claude-3.5 and DeepSeek-R1 excel with a CC Rate of 103.7% and 103.2% separately, while Llama-3.1-8B is notably lower at 53.6%. This difference arises because Claude-3.5 and DeepSeek-R1 often uses concise built-in functions in place of loops. For instance, as illustrated in Fig. 12, when tasked with finding the first element in an array that meets a condition, they employ the target language’s any method from the iterator library, whereas other LLMs rely on traditional loops. Furthermore, the CC Rate of Claude-3.5 and DeepSeek-R1 exceeding 1 indicates that, in terms of cyclomatic complexity, their generated code is simpler than the reference code. Fig. 12 shows an example with C as the source language. This simplification likely results from Rust ground truth developers using the C2Rust migration tool [66], which translates code through one-to-one pattern matching. It is also important to note that although DeepSeek-V3 outperforms Claude-3.5 in terms of pass@1, its CC Rate is 93.5%, which is significantly lower than that of Claude-3.5. This suggests that a model with a lower pass@1 score does not necessarily indicate weaker code capabilities, and the quality of the code it generates may actually be higher.
@@
These findings suggest that DeepSeek-R1 demonstrates a stronger code comprehension and implementation ability, accurately understanding the source code’s functionality and translating it concisely based on the target language’s features.
Figure 11: The performance of LLMs on Key Abilities of code translation.
@@
Figure 12: Claude-3.5 and DeepSeek-R1 translation result compared to Other LLMs and reference code
IV-E3 Conclusion
@@
The complexity of code generated by LLMs reflects their translation capabilities: models with stronger abilities produce code with lower complexity for equivalent functionality. DeepSeek-R1, which excels on RustRepoTrans, also shows superior performance in Noise Robustness, Syntactical Differences Identification, and Code Simplicity. Meanwhile, although Claude-3.5 does not achieve the highest pass@1, it demonstrates exceptional performance in Code Simplicity, highlighting the high quality of the code it generates.
V Discussion
In this section, we further discuss the intended audience and the limitation of RustRepoTrans.
Intended Audience. The RustRepoTrans is designed to provide value to both researchers and practitioners in the field of code translation. Not only do the error causes revealed by RustRepoTrans for different LLMs, along with the associated analysis, offer fine-grained insights for researchers in future work, but the relative performance ranking of various LLMs can also guide developers in choosing suitable base models when constructing their translation tools in real-world scenarios.
Limitations. In our evaluation, the provided information contains comprehensive and non-redundant contextual information. We did not conduct in-depth assessments of two specific aspects: the migration of global variables, type definitions, or function signatures, and the integration of LLMs with static/dynamic analysis. This is because the primary focus of this paper is to achieve a more authentic, practical, and fine-grained evaluation for the precise diagnosis and targeted improvement of LLMs’ code translation capabilities under an incremental translation scenario. However, we provide the complete target codebase to enable any related future work incorporating such analyses.
VI Threats to Validity
One potential threat is data leakage between our benchmark and model training data. However, the training data comprises independent function code from different languages rather than functionally equivalent code pairs, which is crucial for code translation. Another concern is the limited size and variety of programming languages in our benchmark, which may impact the generalizability of our findings. We plan to extend our benchmark in the future. Lastly, the reliability and completeness of the error causes taxonomy pose another potential threat. We employed open coding to systematically identify and categorize the error causes, adhering to established open coding practices to ensure thoroughness and accuracy. Concurrently, we constructed an iterative refining process, reducing individual biases and ensuring more objective assessments. Finally, we engaged a non-participant to annotate the same error messages based on the obtained error cause categories, achieving a Cohen’s kappa of 0.885, indicating almost perfect agreement.
VII Conclusion
@@
This work makes the first attempt to evaluate LLMs on repository-level context code translation targeting Rust. We first manually construct the first repository-level context code translation benchmark RustRepoTrans and evaluate seven representative LLMs. We find that existing LLMs show much worse performance on incremental repository-level context code translation compared to standalone code translation and exhibit more fine-grained deficiencies compared to end-to-end full repository translation. Besides, when the target language is a low-resource language with multiple syntactic constraints, such as Rust, LLMs struggle to effectively identify the various differences between the source and target languages, as well as understand the features of the target language. Meanwhile, we propose a set of more fine-grained evaluation metrics and an enhanced evaluation framework, enabling a more comprehensive analysis of LLMs’ performance in repository-level context code translation tasks to provide fine-grained insights that can effectively inform the development of code translation techniques.
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China (No. 62032025, 62402113), CCF - Sangfor ’Yuanwang’ Research Fund.